### v2

# Contextual Bandit with Herding Effects: Algorithms and Recommendation Applications

Abstract
Contextual bandits serve as a fundamental algorithmic framework for optimizing recommendation decisions online. Though extensive attention has been paid to tailoring contextual bandits for recommendation applications, the "herding effects" in user feedback have been ignored. These herding effects bias user feedback toward historical ratings, breaking down the assumption of unbiased feedback inherent in contextual bandits. This paper develops a novel variant of the contextual bandit that is tailored to address the feedback bias caused by the herding effects. A user feedback model is formulated to capture this feedback bias. We design the TS-Conf (Thompson Sampling under Conformity) algorithm, which employs posterior sampling to balance the exploration and exploitation tradeoff. We prove an upper bound for the regret of the algorithm, revealing the impact of herding effects on learning speed. Extensive experiments on datasets demonstrate that TS-Conf outperforms four benchmark algorithms. Analysis reveals that TS-Conf effectively mitigates the negative impact of herding effects, resulting in faster learning and improved recommendation accuracy.
Keywords:
recommendation contextual bandits herding effects.
1 Introduction
Contextual linear bandit is an important sequential decision making framework for information retrieval applications [5]. It is also applied to optimize news recommendations [10, 20], movie recommendations [16, 17], advertising [32, 22], etc. Recently, a number of variants of contextual linear bandits were proposed to capture important factors of information retrieval applications. Such as the conversational contextual bandit which captures the contextual linear bandit to capture conversational feedbacks in recommendation applications [33], the impatient contextual bandits which captures feedback delay in recommendation applications [13], contextual budgeting bandit which captures the multi-agent nature the budget allocation in online advertising [6], etc.
This paper tackles a critical challenge in the field of recommendation systems: the herding effects in user feedback [1, 28, 24]. Randomized controlled experiments [1, 15] proved the existence of herding effects. The herding effects states that users are conformed to historical ratings of a product when they are assigning ratings. We model the valuation or true preference of a user toward an item as a product of the item’s feature vector and the user’s preference vector. This valuation is unobservable, but a linear combination of the valuation and the historical rating is observable. The weights of this linear combination captures the strength of herding effects serves as a confounder and it is unobservable. This paper presents the first attempt to capture herding effects in contextual linear bandit, and we aim to reveal fundamental understandings on the impact of this unobservable confounder on balancing the exploration vs. exploitation tradeoff for the online recommendation task. Our contribution is the following.
  • We propose a model to quantify herding effects, where a user’s feedback is influenced by both its inherent reward and historical feedback.
  • We develop the TS-Conf algorithm, utilizing Thompson Sampling to balance exploration and exploitation effectively. We provide a regret upper bound for the algorithm, highlighting how herding effects, modulated by the conformity factor, affect learning efficiency.
  • Extensive experiments on four public datasets demonstrate the sublinear regret of the proposed TS-Conf algorithm, and its superior performance over three baselines.
2 Related Work
Contextual linear bandits serve as a fundamental sequential decision making framework for information retrieval applications advertising, recommendation, etc [5]. A number of variants of contextual linear bandits were proposed to capture important factors of information retrieval applications [25, 26, 27, 35]. Conversational contextual bandit tailors the contextual linear bandit to capture conversational feedbacks in recommendation applications [33]. The contextual budgeting bandit extend the contextual linear bandit to the multi-agent for the purpose of studying the budget allocation in online advertising [6]. The key difference to the above work is that our model captures the well-known herding effects in feedback. The new technical challenge is that this herding effects leads to confounded feedback with spurious correlation.
Through controlled experiments[1, 15, 19], some researchers identified a rating bias influenced by historical ratings, observing that users tend to give higher ratings after being exposed to higher historical ratings, a behavior termed as herding effects. Wang et al. [24] introduced an additive generative-based model designed to quantify herding effects. While it can capture the pattern of herding effects, it lacks the neatness required for analytical studies of evolving dynamics of aggregate ratings under herding effects. Krishnan et al. [8] developed a polynomial regression-based model to quantify herding effects. Xie et al. [28] proposed a neater linear model for herding effects and supported analytical studies on the evolving dynamics of aggregate ratings. Other notable psychological effects that lead to biased feedback include assimilate and contrast effects [30], persuasion effects [29], etc. These works focus on rating prediction and are built on the matrix factorization framework. Unlike them, we consider the online decision setting built on the contextual bandit.
A number of works studied bandit learning with biased feedback. Bareinboim et al. [3] extended the reward model of multi-armed bandits such that an unobserved confounder influences the reward. Kallus et al. [7] developed the reward model of multi-armed bandits such that an unobserved instrumental variable influences the reward. Different from them, our work is built on the contextual bandit. Maniu et al. [12] extended the linear bandits to capture social influence bias. In particular, they consider the setting that the preference vectors of users are influenced by their friends and evolve over time. Difference from their work, we apply the confounder model to capture herding effects. Tennenholtz et al. [23] studied linear contextual bandits with access to a large, confounded, offline dataset sampled from some fixed policy. Unlike their work, we consider the setting where an unobserved confounder influences the online reward. Sen et al. [21] studied stochastic contextual bandits with a latent low-dimensional confounder. The confounder is discrete and models the mood of users. Unlike them, our work considers a continuous confounder and the confounder models the strength of herding effects.
3 Model
3.1 The Sequential Decision Framework
We consider the sequential decision problem as one where the decision-maker makes decisions over a finite number of $T\in\mathbb{N_{+}}$ rounds. The set of actions used in the decision-making process is fixed to be a finite set $\mathcal{A}\subset\mathbb{N_{+}}$, where $|\mathcal{A}|<\infty$. Consider a scenario where the decision-maker acts as a movie recommendation system, and $\mathcal{A}$ is the set of movies under consideration. In each round $t\in[T]\triangleq\{{1,...,T}\}$, the decision-maker is presented with a finite set of choices $\mathcal{A}_{t}\subseteq\mathcal{A}$ and $|\mathcal{A}_{t}|=K$, from which it must choose one action ${A_{t}}\in[K]\triangleq\{{1,...,K}\}$ to the user. The user then receives the expected preference reward $\mathbb{E}[R_{t}(A_{t})]$ (positive or negative preference) for the recommending action, which is unobservable to the decision-maker. Based on the reward $\mathbb{E}[R_{t}(A_{t})]$, each user provides feedback $V_{t}(A_{t})\in\mathcal{R}$ about the action $A_{t}$ to decision-maker, where $\mathcal{R}\subset\mathbb{R}$. The application defines the metric for quantifying $V_{t}(A_{t})$. In movie recommendation applications, $V_{t}(A_{t})$ models the user’s rating of the movie.
3.2 The User Feedback Model
Contextual reward model. In our study, we focus on the use of contextual features to evaluate the rewards users receive from recommended items. For each action $a\in\mathcal{A}$, there is a feature vector $\bm{x}_{a}\in\mathbb{R}^{d}$ associated with it that captures contextual information between the user and the action with $d\in\mathbb{N}_{+}$. The preference vector of the user, linked to $\bm{x}_{a}$, is represented as $\bm{\theta}\in\mathbb{R}^{d}$. It should be emphasized that the decision maker has information about the observed context $\bm{x}_{a}$, $\forall a\in\mathcal{A}$, while $\bm{\theta}$ remains unknown. Given $\bm{x}_{a},\bm{\theta}$, in all round $t$, we consider the expected preference reward of the action $a$ from the user, modeled by:
$\mathbb{E}[R_{t}(a)]={\bm{\theta}^{\mathrm{T}}\bm{x}_{a}}.$ (1)
It is important to note that the decision maker cannot observe the reward at each recommendation round. However, after receiving the recommended action, users provide biased feedback. The goal of the decision maker is to maximize the expected cumulative rewards based on this biased user feedback.
User feedback model. We describe user’s feedback depends on both expected reward formed by the user and the historical feedback of actions. Specifically, in applications where ratings serve as feedback, the historical feedback for item $a$ is determined by the rating content, which is known to the decision maker. In each round $t$, let $h_{t,a}\in\mathbb{R_{+}}$ denote the historical feedback of action $a$. Let $\mathbb{E}[R_{t}(a)]$ represent the user’s expected reward. According to the historical feedback $h_{t,a}$ and the expected reward model $\mathbb{E}[R_{t}(a)]$, we define that at time $t$, the user’s feedback to action $a$ is modeled as:
$V_{t}(a)$ $=\alpha h_{t,a}+(1-\alpha)\mathbb{E}[R_{t}(a)]+\eta_{t,a}=\alpha h_{t,a}+(1-\alpha)\bm{\theta}^{\mathrm{T}}\bm{x}_{a}+\eta_{t,a},$ (2)
where $\alpha\in[0,1]$ denotes the conformity tendency of the user, which can be interpreted as the weights that balance between the historical rating $h_{t,a}$ and the expected reward $\mathbb{E}[R_{t}(a)]$. It should be noted that the decision maker has no idea about the exact value of $\alpha$. ${\eta}_{t,a}\in\mathcal{R}$ represents the stochastic noise caused by environmental. Let $f({\eta},\sigma_{a})$ represent the probability density function of ${\eta}_{t,a}$, where $\sigma_{a}$ stands for the unknown standard deviation of ${\eta}_{t,a}$ to the decision maker. The standard deviation $\sigma_{a}$ is also unknown to the decision maker.
3.3 Sequential Decision Making Model
Based on the feedback model, we use $V_{t}(A_{t})$ to study the unknown parameters of the user feedback and to get the user’s preference estimate. To provide a clearer explanation, we consolidate all unknown parameters within $V_{t}(A_{t})$ into $\bm{\Psi}\triangleq[\bm{\theta},\alpha,\bm{\sigma}]$. Here, $\bm{\sigma}\triangleq[\sigma_{1},...,\sigma_{|\mathcal{A}|}]$ represents the noise standard deviation for each action in the action set. After round $t$, the decision-maker possesses information regarding the decision action $A_{t}$, the associated historical feedback $h_{t,A_{t}}$, the feedback $V_{t}(A_{t})$, and the observed context $\bm{x}_{A_{t}}$. Let $\mathcal{H}_{t}$ represent the decision-making history up to decision round $t$ as $\mathcal{H}_{t}\triangleq\{[A_{1},h_{1,A_{1}},\bm{x}_{A_{1}},V_{1}(A_{1})],\\ \ldots,[A_{t},h_{t,A_{t}},\bm{x}_{A_{t}},V_{t}(A_{t})]\}$. In the $t$-th round of decision-making, the decision-maker must base decisions on the history $\mathcal{H}_{t-1}$ of the preceding $t-1$ rounds. Therefore, we propose using a sequential decision-making algorithm that leverages historical dependencies. Specifically, this algorithm maps the decision history to the current round’s action probability distribution $\mathcal{F}(\mathcal{H}_{t-1})$. The action $A_{t}$ is consequently generated from this distribution, expressed as $A_{t}\sim\mathcal{F}(\mathcal{H}_{t-1})$. If the distribution $\mathcal{F}(\mathcal{H}_{t-1})$ targets a single action without variance, we have a deterministic scenario. To measure the effectiveness of a history-dependent algorithm with the probabilistic model $\mathcal{F}$, we present the regret function below:
$R_{T}(\mathcal{F};\bm{\Psi})\triangleq$ $\sum_{t=1}^{T}\max_{a\in\mathcal{A}_{t}}\mathbb{E}[{R}_{t}(a;\bm{\Psi})]-\mathbb{E}\left[R_{t}\left(A_{t}\right)\mid\bm{\Psi},A_{t}\sim\mathcal{F}\left(\mathcal{H}_{t-1}\right)\right],$ (3)
where $\mathbb{E}[{R}_{t}(a;\bm{\Psi})]$ represents the expected reward for action $a$ under parameter $\bm{\Psi}$. A decision maker might have prior knowledge about the preference vector $\bm{\theta}$, the conformity tendency $\alpha$, and the standard deviation $\bm{\sigma}$. We represent this prior knowledge using prior distributions over the parameters, expressed as $p(\bm{\theta})$, $p(\alpha)$, and $p(\sigma_{a})$ for each $a\in\mathcal{A}$. We specifically consider scenarios where $\bm{\theta}$, $\alpha$, and $\sigma_{a}$ independently arise from their respective prior distributions. This relationship is captured by the equation $p(\bm{\Psi})=p(\bm{\theta})p(\alpha)\prod_{a\in\mathcal{A}}p(\sigma_{a})$. The decision maker’s goal is to design a sequential decision-making algorithm based on historical data that minimizes the regret.
4 Algorithm
4.1 Algorithm Design
In the context of accumulating decisions by round $t$, we articulate the model’s parameters, still to be inferred, as $\bm{\Psi}=[\bm{\theta},\alpha,\bm{\sigma}]$. The calculation for their posterior distribution, denoted $p\left(\bm{\Psi}\mid\mathcal{H}_{t}\right)$, is delineated in an ensuing lemma.
Lemma 1. Suppose the probability density function of the noise has the parametric form $f(\cdot,\sigma)$, where $\sigma$ controls the tail property. Given the decision history $\mathcal{H}_{t}$ up to round $t$, the posterior distribution $p\left(\bm{\Psi}\mid\mathcal{H}_{t}\right)$ can be derived as:
$p\left(\bm{\Psi}\mid\mathcal{H}_{t}\right)=$ $\frac{p(\bm{\Psi})}{C}\times\left[\prod\nolimits_{\tau=1}^{t-1}\prod\nolimits_{a\in\mathcal{A}_{\tau}}\left[f\left(\eta_{\tau,a},\sigma_{a}\right)\right]^{\mathbbm{1}_{\left\{A_{\tau}=a\right\}}}\right],$ (4)
$\eta_{\tau,a}=V_{\tau}(a)-\alpha h_{\tau,a}-(1-\alpha)\bm{\theta}^{T}\bm{x}_{a},$ (5)
where $C$ represents the normalizing factor which is independent of the unknown model parameters $\bm{\theta},\alpha,\bm{\sigma}$.
Drawing on the foundational lemma presented earlier, Algorithm 1 introduces a method for posterior sampling tailored to address the challenges of the contextual bandit learning dilemma as discussed in Section 3. Each interaction cycle, or round $t$, commences with the identification of the model’s parameters $\bm{\Psi}$, grounded in the posterior distribution outlined in Eq.(4). Subsequently, the algorithm computes the expected reward for each viable action, with a preference for selecting the action projected to offer the maximum return. Upon the decision maker’s implementation of the chosen action, the system garners feedback $V_{t}(A_{t})$ from the agent. This feedback is subsequently integrated into the decision history, thereby facilitating the transition to the subsequent iteration.
Algorithm 1 TS-Conf (Thompson Sampling under Conformity)
1:  Initialize $\mathcal{H}_{0}=\emptyset$
2:  for $t=1,2,3,\ldots,T$ do
3:     $\bm{\Psi}_{t}\sim p(\bm{\Psi}|\mathcal{H}_{t-1})$ derived in Eq. (4)
4:     Select action by $A_{t}\leftarrow\arg\max_{a\in\mathcal{A}_{t}}\mathbb{E}[{R}_{t}(a;\bm{\Psi}_{t})]$
5:     Observe the user feedback $V_{t}(A_{t})$
6:     Update history $\mathcal{H}_{t}\leftarrow\mathcal{H}_{t-1}\cup[A_{t},h_{t,{A_{t}}},\bm{x}_{A_{t}},V_{t}(A_{t})]$
7:  end for
In general, the posterior distribution derived in Eq. (4) is computationally expensive to sample exactly. Algorithm 2 outlines the design of the TS-ConfMCMC algorithm, which applies the Markov Chain Monte Carlo (MCMC) technique to sample the posterior approximately. Specifically, it is a three-stage Gibbs sampler. It is important to note that Eq.(4) implies that the conditional posterior distribution of ${\sigma}_{a}$ is independent across different actions $a$, given $\bm{\theta}$, $\alpha$. The preference vector $\bm{\theta}$ and the conformity tendency $\alpha$ manifest linearly when two other parameters are provided. We design a three-stage Gibbs sampler to efficiently facilitate the sampling process delineated in step 3 of the TS-Conf algorithm.
Algorithm 2 TS-ConfMCMC
1:  Initialize $\mathcal{H}_{0}=\emptyset$; $\sigma_{a,0}$; $\bm{\theta}_{0}$; ${\alpha}_{0}$
2:  for $t=1,2,3,\ldots,T$ do
3:     for $n=1,2,3,\ldots,N$ do
4:        $\bm{\theta}^{(n)}\sim p(\bm{\theta}|\alpha=\alpha^{(n-1)},\bm{\sigma}=\bm{\sigma}^{(n-1)},\mathcal{H}_{t-1})$
5:        $\alpha^{(n)}\sim p(\alpha|\bm{\sigma}^{(n-1)},\bm{\theta}=\bm{\theta}^{(n)},\mathcal{H}_{t-1})$
6:        $\bm{\sigma}_{a}^{(n)}\sim p(\bm{\sigma}_{a}|\alpha=\alpha^{(n)},\bm{\theta}=\bm{\theta}^{(n)},\mathcal{H}_{t-1}),\forall a$
7:     end for
8:     $\bm{\Psi}_{t}\leftarrow(\bm{\theta}^{(N)},\alpha^{(N)},\bm{\sigma}^{(N)})$
9:     Step 4-6 of Algorithm 1.
10:  end for
4.2 Regret Analysis
Given a parameter $\bm{\Psi}$, denote the estimator for estimating $\alpha$ from $\mathcal{H}_{t}$ as $\widehat{\alpha}_{t}(\bm{\Psi})$. Note that this estimator is defined to assist the proof of regret, we do not need to know how to construct it. We define the confidence bound $\widehat{\alpha}_{t}(\bm{\Psi})$ as:
$\mathbb{P}[\forall t,|\widehat{\alpha}_{t}(\bm{\Psi})-\alpha|\leq W_{t}(\delta;\bm{\Psi})]\geq 1-\delta.$
Different instances of $\widehat{\alpha}_{t}(\bm{\Psi})$ have different confidence width. Let $W^{\ast}_{t}(\delta;\bm{\Psi})$ denote the smallest possible confidence width attained by the optimal estimator $\widehat{\alpha}^{\ast}_{t}(\bm{\Psi})$.
Theorem 4.1
The regret of Algorithm 1 satisfies:
$R_{T}(\mathcal{D})$ $\leq O\left(\frac{1}{1-\alpha}\int\sum\nolimits^{T}_{t=1}W^{\ast}_{t}(1/T;\bm{\Psi})d\bm{\Psi}\right.+\left.\frac{1}{1-\alpha}d\sqrt{T}\ln T\right).$
The following theorem states that the above bound is tight and reveals that our algorithm is guaranteed to converge at a sublinear rate.
Theorem 4.2
If $\sum^{T}_{t=1}W^{\ast}_{t}(1/T;\bm{\Psi})=\Omega(T)$ holds for all $\bm{\Psi}$, the regret is lower bounded by:
$R_{T}(\mathcal{D})\geq\Omega(T).$
The above theorem shows that the algorithm can effectively learn from historical decisions and gradually approach the performance of the best possible decision. Due to page limitations, more details on the proofs of the above two theorems can be found in our technical report [11].
5 Experiments
5.1 Experiment Settings
Datasets. Our approach to constructing the simulated datasets is aligned with established practices in the field, similar to those employed in related studies[31, 14]. In our empirical evaluation using real-world data, we employ datasets sourced from four distinct platforms: Amazon Music [^1] http://jmcauley.ucsd.edu/data/amazon/index\_2014.html, MovieLens [^2] https://grouplens.org/datasets/movielens/, Yelp [^3] https://www.yelp.com/dataset, and Google Maps [^4] http://jmcauley.ucsd.edu/data/googlelocal/. Through the utilization of these varied datasets spanning multiple domains, our objective is to rigorously assess and understand the real-world applicability and performance of the proposed algorithm. Based on the approach similar to [31, 34], we perform data preprocessing and assess the accuracy of algorithm recommendations through regret values. Due to page limitations, more details on the data preprocessing can be found in our technical report [11].
Baselines and metrics. To the best of our knowledge, limited research exists on contextual bandit algorithms that specifically address the herding effects. In light of this gap, we adapt mainstream bandit algorithms to incorporate the herding effects, thereby producing a suitable comparison algorithm. To ensure a fair and relevant comparison in our study, which focuses on the herding effects, we benchmark our proposed algorithm against two sets of algorithms. The first set includes established baseline algorithms, LinUCB[4] and Thompson Sampling (TS)[2], known for their accuracy in scenarios with unbiased user feedback. However, these algorithms do not specifically address herding effects, a gap in the existing research. Recognizing this limitation, we developed LinUCBConf, an adaptation of the LinUCB algorithm. LinUCBConf is designed to provide a more appropriate baseline for our study by accounting for herding effects, which were not explicitly considered in previous models. This adaptation allows for a more equitable comparison, enabling us to effectively demonstrate the strengths and innovations of our proposed algorithm in the context of herding effects. When estimating the preference parameter $\bm{\theta}_{t}$, LinUCBConf employs the same action selection method as LinUCB, using $\widetilde{\bm{x}}_{A(\tau)}$ as the feature vector and $V_{t}(A_{t})$ for feedback to learn user preferences $\bm{\theta}$. To evaluate the efficacy of each algorithm, our primary metric is the regret value, as delineated in Eq.3.
5.2 Stability Analysis of Algorithm Parameters
Impact of the MCMC Approximation. In our study, we focus on analyzing the influence of the number of iterations, denoted as $N$, on the regret metric of our algorithm. The parameter $N$ is critical as it represents the iterations necessary for the MCMC method to approximate samples for the posterior distribution. For the purpose of this analysis, we standardize the dimensions of the actions and the noise variance at $d=10$ and $\sigma^{2}=1.0$, respectively. This uniformity allows for a controlled assessment of the impact of $N$ on the algorithm’s regret. As illustrated in Figure 1, the regret values produced by TS-Conf consistently fall within the range spanned by TS-ConfMCMC. This shows that the approximate algorithm TS-ConfMCMC we proposed can obtain results similar to precise sampling through a limited number of samplings, illustrating the effectiveness of the approximate algorithm.
(a) Dimension $d=5$
(b) Dimension $d=10$
(c) Dimension $d=15$
(d) Dimension $d=20$
(e) Noise Var $\sigma^{2}=0.5$
(f) Noise Var $\sigma^{2}=1.0$
(g) Noise Var $\sigma^{2}=1.5$
(h) Noise Var $\sigma^{2}=2.0$
Figure 1: Comparative analysis of TS-Conf and TS-ConfMCMC
Impact of dimensions $d$. This section delves into the effect of varying dimensionality on the regret value $\hat{R}_{t}$ across different algorithms. We keep the constant noise variance $\sigma^{2}=1.0$ for consistency in our experiments. We evaluate the performance under four dimensions: $d=5$, $d=10$, $d=15$, and $d=20$. Figure 2(a) shows that when the feature dimension is $d=5$, the regret curve for TS-Conf is the lowest among the four algorithms. It suggests that TS-Conf consistently yields lower cumulative regret values than the other three baselines. The LinUCB and TS algorithms exhibit significantly higher regret values than TS-Conf. This observation underscores the pitfalls of assuming unbiased user feedback in the presence of the herding effects, highlighting the necessity to address such biases. While LinUCBConf does account for biased user feedback, its regret values remain higher than those of TS-Conf. This difference underscores the varying outcomes that different exploration-exploitation trade-off strategies can produce. It attests to the efficacy of Bayesian-based posterior sampling techniques in managing uncertainty. As the feature dimension increases to $d=10$, $d=15$, and $d=20$, Figures 2(b), 2(c), and 2(d) mirror the above trends, indicating that these observations persist across different feature dimensions.
(a) Dimension $d=5$
(b) Dimension $d=10$
(c) Dimension $d=15$
(d) Dimension $d=20$
Figure 2: Impact of dimensions $d$ in synthetic dataset
Impact of noise variance $\sigma^{2}$. In this section, we delve into the effect of varying noise variance on the regret value $\hat{R}_{t}$ across different algorithms. Consistently, our experiments are conducted with a feature dimension $d=10$. We assess performance under four distinct noise variance settings: $\sigma^{2}=0.5$, $\sigma^{2}=1.0$, $\sigma^{2}=1.5$, and $\sigma^{2}=2.0$. Figure 3(a) shows that with the noise variance of $\sigma^{2}=0.5$, TS-Conf consistently exhibits the lowest regret values among the four algorithms. LinUCB and TS, which overlook the bias in user feedback, and LinUCBConf, which employs the LinUCB approach for exploration-exploitation trade-off, register significantly higher regret values than TS-Conf. This trend persists as the noise variance increases to $\sigma^{2}=1.0$, $\sigma^{2}=1.5$, and $\sigma^{2}=2.0$, as evidenced in Figure 3(b), Figure 3(c), and Figure 3(d). Such consistent performance under varying noise levels underscores the stability and robustness of the TS-Conf algorithm in handling uncertainties.
(a) Noise Var $\sigma^{2}=0.5$
(b) Noise Var $\sigma^{2}=1.0$
(c) Noise Var $\sigma^{2}=1.5$
(d) Noise Var $\sigma^{2}=2.0$
Figure 3: Impact of noise variance $\sigma^{2}$ in synthetic dataset
5.3 Real-world applications
Results in MovieLens. Figure 4 shows the regret $\hat{R}_{t}$ produced by each algorithm in different dimensions and different noise variances. It can be observed that the TS-Conf algorithm always has the lowest regret value across varying dimensions and noise levels. Consistent with the experimental observations on synthetic data, the regret values for both LinUCB and TS exhibit significantly higher regret than the TS-Conf algorithm and will increase linearly with $t$. Similarly, LinUCBConf also has a greater regret compared with TS-Conf, and its regret values converge at a slower rate than our algorithm.
(a) Dimension $d=5$
(b) Dimension $d=10$
(c) Dimension $d=15$
(d) Dimension $d=20$
(e) Noise Var $\sigma^{2}=0.5$
(f) Noise Var $\sigma^{2}=1.0$
(g) Noise Var $\sigma^{2}=1.5$
(h) Noise Var $\sigma^{2}=2.0$
Figure 4: Impact of dimensions $d$ and noise variance $\sigma^{2}$ in MovieLens dataset.
Results in Yelp. Figure 5 shows the comparison results of the four algorithms on the Yelp dataset. It is evident that the TS-Conf algorithm consistently has the lowest regret value across different dimensions and noise levels. Unlike the LinUCB and TS algorithms, which consistently exhibit linear regret growth, the regret of TS-Conf gradually converges over time. Furthermore, its convergence speed is greater than that of LinUCBConf. Similar results for Amazon Music and Google Maps datasets are presented in our technical report[11].
(a) Dimension $d=5$
(b) Dimension $d=10$
(c) Dimension $d=15$
(d) Dimension $d=20$
(e) Noise Var $\sigma^{2}=0.5$
(f) Noise Var $\sigma^{2}=1.0$
(g) Noise Var $\sigma^{2}=1.5$
(h) Noise Var $\sigma^{2}=2.0$
Figure 5: Impact of dimensions $d$ and noise variance $\sigma^{2}$ in Yelp dataset.
6 Conclusion
This paper presents a contextual bandit framework to address the herding effects problem in recommendation applications. In this framework, we assume that the user feedback on the action is biased and influenced by user preferences and the historical ratings of this action. We design a TS-Conf algorithm that leverages a posterior sampling technique to effectively balance the trade-off between exploration and exploitation. Our theoretical analysis established a sublinear regret bound for the TS-Conf algorithm, demonstrating its efficiency. Extensive experiments on synthetic and real-world datasets show that TS-Conf consistently outperforms three benchmark algorithms, confirming its robustness and superior performance in handling herding effect-induced biases.
Appendix
Proof of Theorem 4.1: Given an arm $a$, one can have:
$(1-\alpha)|\bm{x}^{T}\bm{\theta}-\bm{x}^{T}_{a}\widehat{\bm{\theta}}_{t}|$
$=\left|\begin{array}[]{c c}[\bm{x}^{T}_{a}&0]\end{array}\left[\begin{array}[]{c}(1-\alpha)(\bm{\theta}-\widehat{\bm{\theta}}_{t})\\ 0\end{array}\right]\right|$
$=\left|\begin{array}[]{c c}[\bm{x}^{T}_{a}&0]\end{array}\left[\begin{array}[]{c}(1-\alpha)\bm{\theta}-(1-\alpha)\widehat{\bm{\theta}}_{t}\\ \alpha-\widehat{\alpha}_{t}\end{array}\right]\right|$
$=\left|(\begin{array}[]{c c}[\bm{x}^{T}_{a}&h_{t,a}]\end{array}-\begin{array}[]{c c}[\bm{0}&h_{t,a}]\end{array})\left[\begin{array}[]{c}(1-\alpha)\bm{\theta}-(1-\alpha)\widehat{\bm{\theta}}_{t}\\ \alpha-\widehat{\alpha}_{t}\end{array}\right]\right|$
$=\left|\begin{array}[]{c c}[\bm{x}^{T}_{a}&h_{t,a}]\end{array}\left[\begin{array}[]{c}(1-\alpha)\bm{\theta}-(1-\alpha)\widehat{\bm{\theta}}_{t}\\ \alpha-\widehat{\alpha}_{t}\end{array}\right]\right|-\left|\begin{array}[]{c c}[\bm{0}&h_{t,a}]\end{array}\left[\begin{array}[]{c}(1-\alpha)\bm{\theta}-(1-\alpha)\widehat{\bm{\theta}}_{t}\\ \alpha-\widehat{\alpha}_{t}\end{array}\right]\right|$
$\leq\left|\begin{array}[]{c c}[\bm{x}^{T}_{a}&h_{t,a}]\end{array}\left[\begin{array}[]{c}(1-\alpha)\bm{\theta}-(1-\alpha)\widehat{\bm{\theta}}_{t}\\ \alpha-\widehat{\alpha}_{t}\end{array}\right]\right|+\left|\begin{array}[]{c c}[\bm{0}&h_{t,a}]\end{array}\left[\begin{array}[]{c}(1-\alpha)\bm{\theta}-(1-\alpha)\widehat{\bm{\theta}}_{t}\\ \alpha-\widehat{\alpha}_{t}\end{array}\right]\right|$
$\leq\left|\begin{array}[]{c c}[\bm{x}^{T}_{a}&h_{t,a}]\end{array}\left[\begin{array}[]{c}(1-\alpha)\bm{\theta}-(1-\alpha)\widehat{\bm{\theta}}_{t}\\ \alpha-\widehat{\alpha}_{t}\end{array}\right]\right|+h_{t,a}|\alpha-\widehat{\alpha}_{t}|.$
Note that the first term corresponds to the linear regression problem of estimating the parameter $[(1-\alpha)\bm{\theta}^{T}\alpha]^{T}$, with observation:
$V_{t}(a)$ $=\left[h_{t,a};\bm{x}^{T}_{a}\right]\left[\begin{array}[]{c}\alpha\\ (1-\alpha)\bm{\theta}\end{array}\right]+\eta_{t}.$
The second term corresponding to estimating the strength of herding effects $\alpha$. Thus, the upper confidence bound of $(1-\alpha)|\bm{x}^{T}\bm{\theta}-\bm{x}^{T}_{a}\widehat{\bm{\theta}}_{t}|$ can be bounded by the upper confidence bound of the linear regression problem of estimating $[(1-\alpha)\bm{\theta}^{T}\alpha]^{T}$ plus the upper confidence bound of estimating $\alpha$. Apply [18] and [9], the Bayesian regret can be bounded as
$R^{Bay}_{T}(\mathcal{D})$ $\leq O\left(\frac{1}{1-\alpha}\int\sum^{T}_{t=1}W^{\ast}_{t}(1/T;\bm{\Psi})d\bm{\Psi}\right.+\left.\frac{1}{1-\alpha}d\sqrt{T}\ln T\right).$
This proof is then complete.
Proof of Theorem 4.2: Note that $\alpha\in[0,1]$, implying that $W^{\ast}_{t}(1/T;\bm{\Psi})\leq 1$. The case $\sum^{T}_{t=1}W^{\ast}_{t}(1/T;\bm{\Psi})=\Omega(T)$ implies that there exists a constant $c$ such that $|\alpha-\widehat{\alpha}_{t}|\geq c$ hold for $\Omega(T)$ rounds for any $\bm{\Psi}$. One can select instances of $\bm{\Psi}$ such that the gap between the optimal arm and the sub-optimal arm with the largest the reward is larger or equal to $c$. In such instances, the optimal arm the sub-optimal arm is indistinguishable. Resulting a linear regret of $\Omega(T)$. Defining the prior distribution over such instances, leads to a Bayesian regret of $\Omega(T)$.
A special case. We consider an important special case where exact sampling from the posterior can be computationally efficient. Equation (1) can be rewritten as follows: $V_{t}(A_{t})=\alpha h_{t,{A_{t}}}+(1-\alpha)\bm{\theta}^{\mathrm{T}}\bm{x}_{A(t)}+\eta_{t}=\widetilde{\bm{x}}_{A(t)}^{\text{T}}\widetilde{\bm{\theta}}+\eta_{t,A_{t}}$ where $\widetilde{\bm{\theta}}=\left[\begin{array}[]{c}\alpha\\ (1-\alpha)\bm{\theta}\end{array}\right]$ and $\widetilde{\bm{x}}_{A(t)}=\left[h_{t,{A_{t}}};\bm{x}_{A(t)}\right]$. The special case is composed of Gaussian distributions: (1) the priors of the $\widetilde{\bm{\theta}}$ follows multivariate Gaussian distribution with mean $\bm{\mu}$ and covariance $\bm{\Lambda^{-1}}$, i.e. $p(\widetilde{\bm{\theta}})\sim\mathcal{N}(\bm{\mu},\bm{\Lambda^{-1}})$; (2) the noise $\eta_{t}$ follows a Gaussian distribution, i.e., $f(\eta,\sigma_{a})$ is the density function of $\mathcal{N}(0,\sigma^{2}_{a})$; (3) the variance $\sigma^{2}_{a}$ is given. Under this special case, the posterior $p\left(\widetilde{\bm{\theta}}|\mathcal{H}_{t}\right)$ follows $\mathcal{N}(\bm{\mu_{t}},\bm{\Sigma_{t}})$, where $\bm{\Sigma_{t}}=\Bigl{(}\bm{\Lambda}+\frac{1}{\sigma_{n}^{2}}{\textstyle\sum_{\tau=1}^{t-1}}\widetilde{\bm{x}}_{A(\tau)}\cdot(\widetilde{\bm{x}}_{A(\tau)})^{\mathrm{T}}\Bigr{)}^{-1},$ $\bm{\mu_{t}}=\bm{\Sigma_{t}}(\frac{1}{\sigma_{n}^{2}}{\textstyle\sum_{\tau=1}^{t-1}}V_{\tau}{(A_{\tau})}\cdot\widetilde{\bm{x}}_{A(\tau)}+\bm{\Lambda}\bm{\mu}).$ These formulas imply that sampling the posterior $p\left(\widetilde{\bm{\theta}}|\mathcal{H}_{t}\right)$ is sampling from Gaussian distributions. A sample of $\bm{\theta}$ can be obtained from a sample of $\widetilde{\bm{\theta}}$ as follows: $\bm{\theta}=\widetilde{\bm{\theta}}_{[2:d+1]}/(1-\widetilde{\theta}_{1})$, where $\widetilde{\bm{\theta}}_{[2:d+1]}$ denotes a vector composed of the second to the $(d+1)$-th entries of $\widetilde{\bm{\theta}}$ and $\widetilde{\theta}_{1}$ denotes the first entry of $\widetilde{\bm{\theta}}$.
Implementation details. For synthetic datasets, we generate the observed action features and the user preference vector from Gaussian distributions, represented as $\mathcal{N}(\bm{\mu},\bm{\Sigma})$. Here, $\bm{\mu}$ is a $d$-dimensional mean vector with a value of $\frac{1}{2}$, and $\bm{\Sigma}$ is a $d{\times}d$ covariance matrix with diagonal elements set to $\frac{1}{6}$. And, the user conformity tendency, denoted as $\alpha$, and the historical rating of action, represented as $h_{t,a}$, are assumed to lie within the intervals $[0,1]$ and $[0,5]$, respectively. This approach ensures that our simulations closely resemble real-world scenarios, making them both believable and directly applicable to actual data patterns.
For real-world datasets, the Amazon Music dataset encompasses user ratings for musical items available on Amazon; The MovieLens dataset captures user ratings for movies hosted on the MovieLens platform; The Yelp dataset aggregates user ratings for restaurants listed on Yelp; The Google Maps dataset collates ratings and reviews for various locations and venues available on Google Maps. Table 1 presents the detailed statistics of these datasets.
Table 1: Datasets Summary
Datasets Users Items Ratings
MovieLens 6,040 3,706 1,000,209
Amazon music 478,235 266,414 836,005
Google map 5,054,567 3,116,785 11,453,845
Yelp 366,715 60,785 1,569,264
We conduct our experiments with $T=50000$ decision rounds. In each round, all actions($|\mathcal{A}|=10$) are presented to the decision maker, denoted as $\mathcal{A}_{t}=\mathcal{A}$ for all $t\in[T]$. The noise in the agent’s feedback is simulated according to Eq. (2) using a normal distribution with variance $\sigma^{2}_{a}=1.0$. The density function $f(\eta,\sigma_{a}=1.0)$ corresponds to $\mathcal{N}(\eta,1)$. By default, the feature dimensions are set to $d=10$, unless otherwise specified. The prior distribution of user preference vector follow Gaussian distributions, represented as $\mathcal{N}(\bm{\mu},\bm{\Sigma})$. Here, $\bm{\mu}$ is a $d$-dimensional mean vector with a value of 00, and $\bm{\Sigma}$ is a $d{\times}d$ covariance matrix with diagonal elements set to $1$. The $h_{t,a}$ is set as the average rating of an item. Lastly, prior distribution of $\alpha$ follows $\mathcal{N}(0,1)$.
Data Processing Details of Real-World Datasets. To ensure the integrity of the experiment and mitigate the impact of missing values, we apply certain data filtering criteria. Specifically, we exclude items with a rating count of less than 10 and users with a rating count of less than 10. This filtering process allows us to focus on constructing the necessary dataset for our analysis. To begin with, we calculate the average score for each action $a$ in the dataset, which serves as the historical score $r^{*}_{a}$ for that action. Additionally, we employ the Matrix Factorization (MF) technique to learn item features $\bm{x}_{a}$, user features $\bm{\theta}$, and user conformity tendency $\alpha$. In the case of herding effects, the MF rating model is: $\hat{r}_{ui}=(1-\texttt{Sig}(\beta))\bm{\theta}_{u}^{T}\cdot\bm{x}_{i}+\texttt{Sig}(\beta)r^{*}_{a},$ where $\texttt{Sig}()$ is the sigmoid function, $\texttt{Sig}(\beta)$ represents the estimated value of user conformity tendency $\alpha$. In the model MF, for each user $u$ and item $i$, by comparing the predicted score $\hat{r}^{*}_{ui}$ with the real score $r^{*}_{ui}$ difference to update parameters. We learn the variables $\alpha$, $\bm{x}$ and $\bm{\theta}$ from five dimensions: $d=5$, $d=10$, $d=15$, $d=20$ respectively. We input these inferred variables to the reward model, i.e., Eq. (2) to generate the reward. The noise in the reward follows a normal distribution with variance $\sigma^{2}_{a}=1.0$.
Convergence Analysis. This experiment examines the convergence behavior of the proposed TS-Conf algorithm under varying parameters, notably dimensionality $d$ and noise variance $\sigma^{2}$. Figure 6(a) benchmarks TS-Conf’s convergence across distinct feature dimensions, with $d$ values set to [5, 10, 15, 20], all the while maintaining a consistent noise variance of $\sigma^{2}=1.0$. Conversely, Figure 6(b) evaluates its convergence under a spectrum of noise variances, specifically $\sigma^{2}=[0.5,1.0,1.5,2.0]$, with the dimensionality held constant at $d=10$. A clear pattern emerges: TS-Conf consistently posts the lowest regret values at the minimal settings of $d$ and $\sigma^{2}$. As these parameters escalate, the regret correspondingly surges. This behavior suggests that the algorithm grapples more with the exploration-exploitation trade-off as either $d$ or $\sigma^{2}$ amplifies. This observation aligns our derived regret bound presented in Theorem 4.1. Notably, TS-Conf excels in scenarios with partially observable features or diminished uncertainty levels.
(a) Impact of dimensions $d$
(b) Impact of noise var $\sigma^{2}$
Figure 6: TS-Conf’s performance on different $d$ and $\sigma^{2}$
Results in Google Maps. Figure 7 shows the comparison results of the four algorithms on the Google Maps dataset, respectively. Similarly, it is evident that the TS-Conf algorithm always has the lowest regret value across different dimensions and noise. Differing from the LinUCB and TS algorithms, which consistently exhibit linear regret growth, the regret of TS-Conf gradually converges over time, and the convergence speed is greater than that of LinUCBConf.
(a) Dimension $d=5$
(b) Dimension $d=10$
(c) Dimension $d=15$
(d) Dimension $d=20$
(e) Noise Var $\sigma^{2}=0.5$
(f) Noise Var $\sigma^{2}=1.0$
(g) Noise Var $\sigma^{2}=1.5$
(h) Noise Var $\sigma^{2}=2.0$
Figure 7: Impact of dimensions $d$ and noise variance $\sigma^{2}$ in Google Map dataset.
Results in Amazon Music. Figure 8 shows the regret $\hat{R}_{t}$ produced by each algorithm in different dimensions and different noise variances. It can be observed that the TS-Conf algorithm always has the lowest regret value across varying dimensions and noise levels. In some cases (i.e., $d=15,\sigma^{2}=1.5,\sigma^{2}=2.0$), due to the complexity and uncertainty of real scenarios, the algorithm’s regret value in the initial stages might be higher than the LinUCB and TS algorithms. However, it is notable that both the LinUCB and TS algorithms demonstrate divergence, with their regret values increasing linearly with time ($t$). In contrast, the TS-Conf algorithm shows convergence. Over time, the regret value stabilizes, indicating that the algorithm reaches a steady state of accumulated regret.
(a) Dimension $d=5$
(b) Dimension $d=10$
(c) Dimension $d=15$
(d) Dimension $d=20$
(e) Noise Var $\sigma^{2}=0.5$
(f) Noise Var $\sigma^{2}=1.0$
(g) Noise Var $\sigma^{2}=1.5$
(h) Noise Var $\sigma^{2}=2.0$
Figure 8: Impact of dimensions $d$ and noise variance $\sigma^{2}$ in Amazon dataset.